Circular No. 8227

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 U1 (LINEAR) L. Manguso and H. Stange, Lincoln Laboratory, Massachusetts Insti-

tute of Technology, report the discovery of a comet with a definite halo but no tail on LINEAR images (discovery observation below). Following posting on the NEO Confirmation Page, two other CCD observers have also commented on the cometary appearance: J. E. McGaha (Tucson, AZ, 0.30-m reflector; faint 10′′ coma elongated in p.a. 240^{o} on Oct. 20.3 UT) and J. Young (Table Mountain, CA, 0.6-m reflector; coma of diameter 6′′ and mag 17.5 with a tail ∼ 16′′ long in p.a. 276^{o}).

2003 UT  α_{2000}  δ_{2000}  Mag.
Oct. 19.37824  5^{h}27^{m}58.^{s}79  −3^{o}17′53′′.0  18.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-U25.

T = 2003 Nov. 1.971 TT  ω = 276.^{o}587
Ω = 321.837 }2000.0
q = 1.80838 AU  i = 164.381

1937 UB (HERMES) J. L. Margot, University of California, Los Angeles; M. C. Nolan, V.

Negron, A. A. Hine, D. B. Campbell, and E. S. Howell, National Astronomy and Ionosphere Center; L. A. M. Benner, S. J. Ostro, and J. D. Giorgini, Jet Propulsion Laboratory; and B. G. Marsden, Minor Planet Center, write: “Arecibo (2380-MHz, 12.6-cm) radar observations on Oct. 18 and 20 UT show that the recently rediscovered (IAUC 8223) near-earth-object 1937 UB has a strongly bifurcated appearance. Our images show two separate components of roughly equal sizes, consistent with an orbiting binary pair. The range-Doppler separation between components was 150 m at 1.5 Hz on Oct. 18.2 and 600 m at 0.7 Hz on Oct. 20.1. The Doppler broadening of each component was 0.8 Hz on Oct. 18 and 0.6 Hz on Oct. 20. Preliminary estimates of the diameters, based on visible range extents, are 300–450 m. Upper limits to the spin period of each component are 13–21 hr for the above size range. Additional data are needed to provide an unambiguous orbital solution and to verify whether the system is in a doubly synchronous configuration. Radar and lightcurve observations throughout this apparition would be extremely valuable.”

2003 October 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green